# A Simple, Optimal and Efficient Algorithm for Online Exp-Concave Optimization

### v1

@@ D.3 Proofs for Memory-Efficient OXO
$\leq$ $\frac{1}{2\gamma^{\prime}}\left\|\nabla g_{t}(\mathbf{y}_{t})\right\|_{\widetilde{A}_{t}^{-1}}^{2}+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{\widetilde{A}_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{\widetilde{A}_{t}}^{2}$
where the last inequality uses the fact that $\widetilde{A}_{t}\preceq\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$, which is ensured by Algorithm˜7. Then plugging Lemma˜14 into the preceding inequality yields
$f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})$ $\leq\frac{1}{2\gamma^{\prime}}\left(\left\langle\widetilde{A}_{t}^{-1},\widetilde{A}_{t}-\widetilde{A}_{t-1}\right\rangle_{\operatorname{F}}+\Delta_{t}\right)+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{\widetilde{A}_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{\widetilde{A}_{t}}^{2}$
$\leq\frac{1}{2\gamma^{\prime}}\left(\log\operatorname{det}(\widetilde{A}_{t})-\log\operatorname{det}(\widetilde{A}_{t-1})+\Delta_{t}\right)+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{\widetilde{A}_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{\widetilde{A}_{t}}^{2}$
where the equality uses the fact that $\left\langle X^{-1},X-Y\right\rangle_{\operatorname{F}}\leq\log\operatorname{det}(X)-\log\operatorname{det}(Y)$, which comes from the proof of Lemmas˜5 and 6. Telescoping the preceding inequality and discarding the negative terms establishes the desired regret bound Eq.˜32:
$\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\right)\leq\frac{1}{2\gamma^{\prime}}\log\frac{\operatorname{det}(\widetilde{A}_{T})}{\operatorname{det}(\widetilde{A}_{0})}+\frac{1}{2\gamma^{\prime}}\Delta_{1:T}+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{1}-\mathbf{u}\right\|_{\widetilde{A}_{0}}^{2}.$
The logarithmic term is further bounded with Jensen’s inequality, differing slightly from Lemma˜5 due to the number of non-zero eigenvalues:
@@
Finally, the runtime Eq.˜34 follows from the following two parts:
  • Runtime aside from Algorithm˜7. Following the same analysis of Lemmas˜2 and 13, overall runtime aside from Algorithm˜7 is $O((k-1)^{-1}\sqrt{(d+\Delta_{1:T})T/\epsilon}\cdot d^{\omega}\log T)$ due to FastProj. It suffices to verify the following inequality which follows from Lemma˜15:
@@
  • Runtime of Algorithm˜7. After SVD is performed, the last $d^{\prime}+1$ rows of $S_{t}$ are all-zero, thus the SVD is performed at most $\lceil T/(d^{\prime}+1)\rceil$ times. The runtime of SVD is $O(d^{\prime 2}d\log T)$ [Golub and Van Loan, 2013] to achieve the desired accuracy that does not affect the regret bound.  [^4] The runtime of SVD can be improved to $O\big(d^{\prime 2}(d+\log T)\big)$ and the factor $O(\log T)$ can be independent of the minimal singular value gap of $S_{t}$ and only depends on the scale of singular values of $S_{t}$ [Parlett, 1998]. The runtime of updating $R_{t}$ is $O(d^{\prime}d)$, as this can be implemented with twice rank-one updates:
@@
    $=\underbrace{\epsilon I+S_{t-1}S_{t-1}^{\mathsf{T}}}_{R_{t-1}^{-1}}+\mathbf{e}_{i_{t}}\underbrace{\left(S_{t-1}\nabla g_{t}(\mathbf{y}_{t})\right)^{\mathsf{T}}}_{\mathbf{a}_{t}^{\mathsf{T}}}+\underbrace{\left(S_{t-1}\nabla g_{t}(\mathbf{y}_{t})+\mathbf{e}_{i_{t}}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right)}_{\mathbf{b}_{t}}\mathbf{e}_{i_{t}}^{\mathsf{T}}$
    where $\mathbf{e}_{i_{t}}$, $\mathbf{a}_{t}$ and $\mathbf{b}_{t}$ are $2d^{\prime}$-dimensional vectors. Therefore, the overall runtime of Algorithm˜7 is $O(d^{\prime 2}d(\log T)\lceil T/(d^{\prime}+1)\rceil+d^{\prime}dT)=O(d^{\prime}dT\log T)$.

### v2

@@ E.3 Proofs and Details for Memory-Efficient OXO
[…]
where the last inequality uses the fact that $\widetilde{A}_{t}\preceq\widetilde{A}_{t-1}+\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}$, which is ensured by Algorithm˜7. Then plugging Lemma˜15 into the preceding inequality yields
$\quad f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})$
$\leq\frac{1}{2\gamma^{\prime}}\left(\left\langle\widetilde{A}_{t}^{-1},\widetilde{A}_{t}-\widetilde{A}_{t-1}\right\rangle_{\operatorname{F}}+\Delta_{t}\right)+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{\widetilde{A}_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{\widetilde{A}_{t}}^{2}$
$\leq\frac{1}{2\gamma^{\prime}}\left(\log\operatorname{det}(\widetilde{A}_{t})-\log\operatorname{det}(\widetilde{A}_{t-1})+\Delta_{t}\right)+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t}-\mathbf{u}\right\|_{\widetilde{A}_{t-1}}^{2}-\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{t+1}-\mathbf{u}\right\|_{\widetilde{A}_{t}}^{2},$
where the equality uses the fact that $\left\langle X^{-1},X-Y\right\rangle_{\operatorname{F}}\leq\log\operatorname{det}(X)-\log\operatorname{det}(Y)$, which comes from the proof of Lemmas˜5 and 6. Telescoping the preceding inequality yields
$\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\right)\leq\frac{1}{2\gamma^{\prime}}\log\frac{\operatorname{det}(\widetilde{A}_{T})}{\operatorname{det}(\widetilde{A}_{0})}+\frac{1}{2\gamma^{\prime}}\Delta_{1:T}+\frac{\gamma^{\prime}}{2}\left\|\mathbf{y}_{1}-\mathbf{u}\right\|_{\widetilde{A}_{0}}^{2},$
where the logarithmic term is further bounded with Jensen’s inequality, which differs slightly from Lemma˜5 due to the number of non-zero eigenvalues:
@@
Finally, the runtime follows from the following two parts:
  • Runtime aside from Algorithm˜7. Following the same analysis of Lemmas˜2 and 14, overall runtime aside from Algorithm˜7 is $O((k-1)^{-1}\sqrt{(d+\Delta_{1:T})T/\epsilon}\cdot d^{\omega}\log T)$ due to FastProj. It suffices to verify the following inequality which follows from Lemma˜16:
@@
  • Runtime of Algorithm˜7. Each time when SVD is triggered, the last $d^{\prime}+1$ rows of $S_{t}$ become all-zero, thus the SVD is triggered at most $\lceil T/(d^{\prime}+1)\rceil$ times. The runtime of SVD is $O(d^{\prime 2}d\log T)$ [Golub and Van Loan, 2013] to achieve the desired accuracy that does not affect the regret bound.  [^5] The runtime of SVD can be improved to $O\big(d^{\prime 2}(d+\log T)\big)$ and the factor $O(\log T)$ can be independent of the minimal singular value gap of $S_{t}$ and only depends on the scale of singular values of $S_{t}$ [Parlett, 1998]. The runtime of updating $R_{t}$ is $O(d^{\prime}d)$, as this can be implemented with twice rank-one updates:
@@
    $=\underbrace{\epsilon I+S_{t-1}S_{t-1}^{\mathsf{T}}}_{R_{t-1}^{-1}}+\mathbf{e}_{i_{t}}\underbrace{\left(S_{t-1}\nabla g_{t}(\mathbf{y}_{t})\right)^{\mathsf{T}}}_{\mathbf{a}_{t}^{\mathsf{T}}}+\underbrace{\left(S_{t-1}\nabla g_{t}(\mathbf{y}_{t})+\mathbf{e}_{i_{t}}\nabla g_{t}(\mathbf{y}_{t})\nabla g_{t}(\mathbf{y}_{t})^{\mathsf{T}}\right)}_{\mathbf{b}_{t}}\mathbf{e}_{i_{t}}^{\mathsf{T}},$
    where $\mathbf{e}_{i_{t}}$, $\mathbf{a}_{t}$ and $\mathbf{b}_{t}$ are $2d^{\prime}$-dimensional vectors that can be computed in $O(d^{\prime}d)$ time. Therefore, the overall runtime of Algorithm˜7 is $O(d^{\prime 2}d(\log T)\lceil T/(d^{\prime}+1)\rceil+d^{\prime}dT)=O(d^{\prime}dT\log T)$.
Comparison with OQNS.
Mhammedi and Gatmiry [2023] mention the possibility of combining OQNS with sketching but provide neither algorithms nor analysis. Incorporating sketching further complicates OQNS’s already intricate analysis, as the sketching error interacts with the log-barrier, Hessian approximation and decisions’ update in Eq.˜4. In contrast, within the OMD framework, sketching errors naturally appear as an additive term in the regret bound via OMD’s stability term, illustrated in Eq.˜29, since OMD’s regret decomposition is somewhat orthogonal to sketching.